Calculate the number of grams-formula weights in 100 g of hydrated copper sulfate


1
Expert's answer
2021-03-01T02:26:01-0500

n = m/M

M (Cu(HSO4)2) = 257.69 g/mol

n (Cu(HSO4)2) = 100/257.69 = 0.39 mol